# Consistency of Neural Causal Partial Identification

### v1

@@ C.2 Proof of Theorem 4
    [52, Proposition 1] shows that given a $L$-Lipschitz continuous function $f:[0,1]^{m}\rightarrow\mathbb{R}$, there exist $\hat{f}\in\mathcal{NN}_{m,1}(W,\Theta(\log(m))$ such that $\|f-\hat{f}\|_{\infty}\leqslant O\left(W^{-1/d}\right)$ and $\hat{f}$ is $\sqrt{m}L$-Lipschitz continuous.  [^2] The original proof does not specify the depth of the neural network. The authors show that the network architectures can be chosen as consisting of $O(W)$ parallel blocks each having the same architecture. Each block is used to realize the function $\phi(x)$ $=(\min(\min_{k\neq s}(x_{k}-x_{s}+1),\min_{k}(1+x_{k}),\min_{k}(1-x_{k})))_{+},\quad x\in\mathbb{R}^{m}.$ This function can be realized by feed-forward network with width $O\left(m^{2}\right)$ and depth $O(\log m)$. For a multivariate function with output dimension $m^{\prime}$ , we can approximate each coordinate individually and get a neural network approximation $\hat{f}$ that is $\sqrt{mm^{\prime}}L$-Lipschitz continuous and $\|f-\hat{f}\|_{\infty}\leqslant O\left(W^{-1/m}\right)$.
@@
    $\|f-T_{K}\hat{f}\|_{\infty}$ $=\|T_{K}f-T_{K}\hat{f}\|_{\infty}\leqslant\|f-\hat{f}\|_{\infty}\leqslant O\left(W^{-1/m}\right).$
@@
    $\sup_{f\in\mathcal{F}_{L}^{K}([-K,K]^{m},[-K,K]^{m^{\prime}})}\inf_{\hat{f}\in\mathcal{NN}_{m,m^{\prime}}^{\sqrt{mm^{\prime}}{L},K}\left(W_{0,n},\Theta\left(\log m\right)\right)}\|f-\hat{f}\|_{\infty}\leqslant O\left(W_{0,n}^{-1/m}\right).$ (29)
    Theorem 3 implies that for each latent variable $U_{i}$, there exist a neural network $g_{i}$ with architecture in Corollary 1 such that
@@
    By 3, we know $\|U_{i}\|_{\infty}\leqslant K$, which implies
    $W(P(U_{i}),P((T_{K}g_{i})(Z_{i})))\leqslant W(P(U_{i}),P(g_{i}(Z_{i})))\leqslant O(\sum_{i=1}^{n}L^{-2/d^{U}_{\max}}_{i,n}+\tau_{n}(1-\log\tau_{n})).$ (30)
    Combining Equations 29 and 30 with the same proof as Corollary 1, it can be proven in the same way as Corollary 1 that there exist a $\theta_{n}\in\Theta_{n}$ satisfying
@@
    Let $\tilde{\Theta}_{\infty}=\{\theta^{*}\}$, we have verified the third assumption in Proposition 6.
Take the Wasserstein radius to be $\alpha_{n}=O(s_{n}+\epsilon_{n})$, Proposition 6 implies the conclusion. ∎
C.3 Proof of Proposition 3
Lemma 2.
Let $(\hat{T},\hat{Y})\sim\mu,(T,Y)\sim\nu$ and suppose that $f(t)=\mathbb{E}_{\nu}[Y|T],\hat{f}(t)=\mathbb{E}_{\mu}[\hat{Y}|\hat{T}]$ are $L$-Lipschitz continuous and $|f(t)|\leqslant K,|\hat{f}(t)|\leqslant K$, then we have
@@ C.3 Proof of Proposition 3
$W(\mu,\nu)=\sup_{g\in\text{Lip}(1)}\mathbb{E}_{\mu}[g(\hat{T},\hat{Y})]-\mathbb{E}_{\nu}[g(T,Y)].$
Let $g_{0}(t,y)=(\hat{f}(t)-f(t))y$, we verify $g_{0}$ is a Lipschitz continuous function in $\{(t,y):\|(t,y)\|_{\infty}\leqslant K\}$.
@@
$=|(\hat{f}(t_{1})-f(t_{1}))(y_{1}-y_{2})|+|y_{2}(\hat{f}(t_{1})-f(t_{1})-(\hat{f}(t_{2})-f(t_{2})))|$
$\leqslant 2K|y_{1}-y_{2}|+K(|\hat{f}(t_{1})-\hat{f}(t_{2})|+|f(t_{1})-f(t_{2})|)$
@@
$W(\mu,\nu)$ $\geqslant\frac{1}{L_{g}}(\mathbb{E}_{\mu}[g_{0}(\hat{T},\hat{Y})]-\mathbb{E}_{\nu}[g_{0}(T,Y)])$
$=\frac{1}{L_{g}}(\mathbb{E}_{\mu}[(\hat{f}(\hat{T})-f(\hat{T}))\mathbb{E}[\hat{Y}|\hat{T}=t]]-\mathbb{E}_{\nu}[(\hat{f}(T)-f(T))\mathbb{E}[Y|T=t]])$
$=\frac{1}{L_{g}}(\mathbb{E}_{\mu}[(\hat{f}(\hat{T})-f(\hat{T}))\hat{f}(\hat{T})]-\mathbb{E}_{\nu}[(\hat{f}(T)-f(T))f(T)]).$ (31)
Now, let $h(t)=(\hat{f}(t)-f(t))\hat{f}(t)$. Following the same argument, it can be proven that $h(t)$ is Lipschitz continuous with Lipschitz constant being $L_{h}=4LK$. Hence,
$\mathbb{E}_{\mu}[h(\hat{T})]$ $\geqslant\mathbb{E}_{\nu}[h(T)]-L_{h}W(\mu,\nu).$
Plug into (31), and we get
$(L_{g}+L_{h})W(\mu,\nu)$ $\geqslant\mathbb{E}_{\nu}[(\hat{f}(T)-f(T))\hat{f}(T)-(\hat{f}(T)-f(T))f(T)]$
$=\mathbb{E}_{\nu}\left[(\hat{f}(T)-f(T))^{2}\right].$
Proof of Proposition 3.
@@
$=(t_{2}-t_{1})(\mathbb{E}_{{\mathcal{M}}}[Y]-\mathbb{E}_{\hat{\mathcal{M}}}[\hat{Y}])^{2}\leqslant(t_{2}-t_{1})W(P^{\mathcal{M}}(\boldsymbol{V}),P^{\hat{\mathcal{M}}}(\hat{\boldsymbol{V}}))^{2}.$
Now, suppose that $Y$ is a descendant of $T$. Let $X=\text{Pa}(T)$. Note that $f_{y}(x,t)=\mathbb{E}_{{\mathcal{M}}}[Y|X=x,T=t]$ is $L_{y}$-Lipschitz continuous with Lipschitz constant $L_{y}\leqslant(L+1)^{n_{V}}$. This is because $Y$ can be written as $Y=F_{0}(X,T,\boldsymbol{U}_{y})$ where $\boldsymbol{U}_{y}$ are latent variables independent of $T,X$ and $F_{0}$ is composition of $f_{i}$ in structure equations. The composition of Lipschitz functions is still Lipschitz and $F_{0}$ is a composition of at most $n_{V}$ $L$-Lipschitz functions. $F_{0}$ is $(L+1)^{n_{V}}$-Lipschitz and so is $f_{y}(x,t)=\mathbb{E}_{\boldsymbol{U}_{y}}[F_{0}(X,T,\boldsymbol{U}_{y})]$. Recall that $\nu,\mu$ are the observation distributions of $\mathcal{M}$ and $\hat{\mathcal{M}}$ respectively in Proposition 3. Similarly, $\hat{f}_{y}(x,t)=\mathbb{E}_{\hat{\mathcal{M}}}[\hat{Y}|X=x,T=t]$ is $L_{y}$-Lipschitz continuous. By Lemma 2 and the overlap assumption, we have
$C_{W}W(\mu,\nu)$ $\geqslant\mathbb{E}_{\nu}\left[(f_{y}(X,T)-\hat{f}_{y}(X,T))^{2}\right]$
$=\int(f_{y}(x,t)-\hat{f}_{y}(x,t))^{2}\nu(dt|x)\nu(dx)$
$\geqslant\delta\int_{t_{1}}^{t_{2}}\int(f_{y}(x,t)-\hat{f}_{y}(x,t))^{2}\nu(dx)P(dt).$
$\geqslant\delta\int_{t_{1}}^{t_{2}}\left(\int f_{y}(x,t)\nu(dx)-\int\hat{f}_{y}(x,t)\nu(dx)\right)^{2}P(dt).$
Note that $\hat{f}_{y}(x,t)$ is Lipschitz continuous, we have
$\Bigl{|}\int\hat{f}_{y}(x,t)\nu(dx)-\int\hat{f}_{y}(x,t)\mu(dx)\Bigl{|}\leqslant L_{y}W(\mu,\nu).$
@@
$\left(\int f_{y}(x,t)\nu(dx)-\int\hat{f}_{y}(x,t)\nu(dx)\right)^{2}$ $\geqslant\frac{1}{2}\left(\int f_{y}(x,t)\nu(dx)-\int\hat{f}_{y}(x,t)\mu(dx)\right)^{2}$
$\ \ \ \ \ \ \ \ \ \ \ \ -\left(\int\hat{f}_{y}(x,t)\mu(dx)-\int\hat{f}_{y}(x,t)\nu(dx)\right)^{2}$
$\geqslant\frac{1}{2}\left(\int f_{y}(x,t)\nu(dx)-\int\hat{f}_{y}(x,t)\mu(dx)\right)^{2}-L_{y}^{2}W^{2}(\mu,\nu)$
$=\frac{1}{2}(\mathbb{E}_{\mathcal{M}}[Y(t)]-\mathbb{E}_{\hat{\mathcal{M}}}[\hat{Y}(t)])^{2}-L_{y}^{2}W^{2}(\mu,\nu).$
@@
$\int_{t_{1}}^{t_{2}}(\mathbb{E}_{\mathcal{M}}[Y(t)]-\mathbb{E}_{\hat{\mathcal{M}}}[\hat{Y}(t)])^{2}P(dt)$ $\leqslant\frac{2C_{W}}{\delta}W(\mu,\nu)+2L_{y}^{2}W^{2}(\mu,\nu)(t_{2}-t_{1}).$
Proof of Corollary 2.
In the proof of Theorem 4, we know with probability at least $1-O(n^{-2})$, (6) has feasible solutions and
$W(P_{n}^{\mathcal{M^{*}}}(\boldsymbol{V}),P^{\mathcal{M^{*}}}(\boldsymbol{V}))\leqslant O(\alpha_{n}),\quad W(P_{m_{n}}^{\theta_{n}}(\boldsymbol{V}),P^{\theta_{n}}(\boldsymbol{V}))\leqslant O(\alpha_{n}),$
where notations $\theta$ and $P^{\theta}$ are defined in the proof of Theorem 4 and $\theta_{n}$ is one of the minimizers of (6). By Lemma 7, we know that
$W(P_{n}^{\mathcal{M^{*}}}(\boldsymbol{V}),P_{m_{n}}^{\theta_{n}}(\boldsymbol{V}))\leqslant S_{\lambda_{n}}(P_{n}^{\mathcal{M^{*}}}(\boldsymbol{V}),P_{m_{n}}^{\theta_{n}}(\boldsymbol{V}))+2(\log(m_{n}n)+1)\lambda_{n}\leqslant O(\alpha_{n}).$
@@
$W(P^{\mathcal{M^{*}}}(\boldsymbol{V}),P^{\theta_{n}}(\boldsymbol{V}))\leqslant W(P_{n}^{\mathcal{M^{*}}}(\boldsymbol{V}),P^{\mathcal{M^{*}}}(\boldsymbol{V}))+W(P^{\mathcal{M^{*}}}(\boldsymbol{V}),P^{\theta_{n}}(\boldsymbol{V}))+W(P_{m_{n}}^{\theta_{n}}(\boldsymbol{V}),P^{\theta_{n}}(\boldsymbol{V}))\leqslant O(\alpha_{n}).$
By Proposition 3, we get $|F_{n}-F_{*}|\leqslant O(\sqrt{\alpha_{n}})$. ∎
@@ Appendix D Experiments
The structure equations of the generative models are (5). We use three-layer feed-forward neural networks with width 128 for each $\hat{f}_{i}$ and six-layer neural networks with width 128 for each $\hat{g}_{j}$. We use the Augmented Lagrangian Multiplier (ALM) method to solve the optimization problems as in [3]. We run $600$ epochs and use a batch size of 2048 in each epoch. $m_{n}$ is set to be $m_{n}=n$. The "geomloss" package [16] is used to calculate the Sinkhorn distance. To impose Lipschitz regularization, we use the technique from [23] to do layer-wise normalization to the weight matrices with respect to infinity norm. The upper bound of the Lipschitz constant in each layer is set to be $8$. The $\tau$ in the Gumbel-softmax layer is set to be 00.
For the choice of Wasserstein ball radius $\alpha_{n}$, we use the subsampling technique from [11, Section 3.4]. We can take the criterion function in [11] to be $Q(\theta)=W(P^{\mathcal{M}^{*}}(\boldsymbol{V}),P^{\mathcal{M}^{\theta}}(\boldsymbol{V}))$ and its empirical estimation to be $Q_{n}(\theta)=W(P_{n}^{\mathcal{M}^{*}}(\boldsymbol{V}),P_{n}^{\mathcal{M}^{\theta}}(\boldsymbol{V}))$. In the first step, we minimize the Wasserstein distance $\hat{\theta}^{*}_{n}=\arg\min_{\theta}W(P^{\mathcal{M}^{*}}_{n}(\boldsymbol{V}),P^{\mathcal{M}^{\theta}}_{n}(\boldsymbol{V}))$. Then, [11] propose to refine the radius $Q_{n}(\hat{\theta}^{*}_{n})$ by taking the quantile of subsample $\{\sup_{\theta:Q_{n}(\hat{\theta})\leqslant\beta Q_{n}(\hat{\theta}^{*}_{n})}Q_{j,b}(\theta)\}$, where $Q_{j,b}$ denotes the criterion function estimated at $j$-th subsample of size $b$. However, it is time-consuming to solve this optimization problem many times. In practice, we set the radius $\alpha_{n}$ to be the $95\%$ quantile of $\{W(\bar{P}_{j}^{\mathcal{M}^{*}}\}(\boldsymbol{V}),P^{\mathcal{M}^{\hat{\theta}^{*}_{n}}}_{n}(\boldsymbol{V}))$, where we use 50 subsamples $\bar{P}_{j}^{\mathcal{M}^{*}},j=1,\cdots,50$ from $P^{\mathcal{M}^{*}_{n}(\boldsymbol{V})}$ with size $b=\lfloor 15\sqrt{n}\rfloor$.

### v2

@@ Appendix C Proof of Consistency
    [53, Proposition 1] shows that given a $L$-Lipschitz continuous function $f:[0,1]^{m}\rightarrow\mathbb{R}$, there exist $\widehat{f}\in\mathcal{NN}_{m,1}(W,\Theta(\log(m))$ such that $\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/d}\right)$ and $\widehat{f}$ is $\sqrt{m}L$-Lipschitz continuous.  [^6] The original proof does not specify the depth of the neural network. The authors show that the network architectures can be chosen as consisting of $O(W)$ parallel blocks each having the same architecture. Each block is used to realize the function $\phi(x)$ $=(\min(\min_{k\neq s}(x_{k}-x_{s}+1),\min_{k}(1+x_{k}),\min_{k}(1-x_{k})))_{+},\quad x\in\mathbb{R}^{m}.$ This function can be realized by feed-forward network with width $O\left(m^{2}\right)$ and depth $O(\log m)$. For a multivariate function with output dimension $m^{\prime}$ , we can approximate each coordinate individually and get a neural network approximation $\widehat{f}$ that is $\sqrt{mm^{\prime}}L$-Lipschitz continuous and $\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/m}\right)$.
@@
    $\|f-T_{K}\widehat{f}\|_{\infty}$ $=\|T_{K}f-T_{K}\widehat{f}\|_{\infty}\leqslant\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/m}\right).$
@@
    $\sup_{f\in\mathcal{F}_{L}^{K}([-K,K]^{m},[-K,K]^{m^{\prime}})}\inf_{\widehat{f}\in\mathcal{NN}_{m,m^{\prime}}^{\sqrt{mm^{\prime}}{L},K}\left(W_{0,n},\Theta\left(\log m\right)\right)}\|f-\widehat{f}\|_{\infty}\leqslant O\left(W_{0,n}^{-1/m}\right).$ (C.9)
    Theorem 3.8 implies that for each latent variable $U_{i}$, there exist a neural network $g_{i}$ with architecture in Corollary 3.9 such that
@@
    By 3.2, we know $\|U_{i}\|_{\infty}\leqslant K$, which implies
    $W(P(U_{i}),P((T_{K}g_{i})(Z_{i})))\leqslant W(P(U_{i}),P(g_{i}(Z_{i})))\leqslant O(\sum_{i=1}^{n}L^{-2/d^{U}_{\max}}_{i,n}+\tau_{n}(1-\log\tau_{n})).$ (C.10)
    Combining Equations C.9 and C.10 with the same proof as Corollary 3.9, it can be proven in the same way as Corollary 3.9 that there exists a $\theta_{n}\in\Theta_{n}$ satisfying
@@
    Let $\widetilde{\Theta}_{\infty}=\{\theta^{*}\}$, we have verified the third assumption in Proposition C.2.
Take the Wasserstein radius to be $\alpha_{n}=O(s_{n}+\epsilon_{n})$, Proposition C.2 implies the conclusion. ∎
C.3 Proof of Proposition 4.3
Lemma C.3.
Let $(\widehat{T},\widehat{Y})\sim\mu,(T,Y)\sim\nu$ and suppose that $f(t)=\mathbb{E}_{\nu}[Y|T],\widehat{f}(t)=\mathbb{E}_{\mu}[\widehat{Y}|\widehat{T}]$ are $L$-Lipschitz continuous and $|f(t)|\leqslant K,|\widehat{f}(t)|\leqslant K$, then we have
@@
$W(\mu,\nu)=\sup_{g\in\text{Lip}(1)}\mathbb{E}_{\mu}[g(\widehat{T},\widehat{Y})]-\mathbb{E}_{\nu}[g(T,Y)].$
Let $g_{0}(t,y)=(\widehat{f}(t)-f(t))y$, we verify $g_{0}$ is a Lipschitz continuous function in $\{(t,y):\|(t,y)\|_{\infty}\leqslant K\}$.
@@
$=|(\widehat{f}(t_{1})-f(t_{1}))(y_{1}-y_{2})|+|y_{2}(\widehat{f}(t_{1})-f(t_{1})-(\widehat{f}(t_{2})-f(t_{2})))|$
$\leqslant 2K|y_{1}-y_{2}|+K(|\widehat{f}(t_{1})-\widehat{f}(t_{2})|+|f(t_{1})-f(t_{2})|)$
@@
$W(\mu,\nu)$ $\geqslant\frac{1}{L_{g}}(\mathbb{E}_{\mu}[g_{0}(\widehat{T},\widehat{Y})]-\mathbb{E}_{\nu}[g_{0}(T,Y)])$
$=\frac{1}{L_{g}}(\mathbb{E}_{\mu}[(\widehat{f}(\widehat{T})-f(\widehat{T}))\mathbb{E}[\widehat{Y}|\widehat{T}=t]]-\mathbb{E}_{\nu}[(\widehat{f}(T)-f(T))\mathbb{E}[Y|T=t]])$
$=\frac{1}{L_{g}}(\mathbb{E}_{\mu}[(\widehat{f}(\widehat{T})-f(\widehat{T}))\widehat{f}(\widehat{T})]-\mathbb{E}_{\nu}[(\widehat{f}(T)-f(T))f(T)]).$ (C.11)
Now, let $h(t)=(\widehat{f}(t)-f(t))\widehat{f}(t)$. Following the same argument, it can be proven that $h(t)$ is Lipschitz continuous with Lipschitz constant being $L_{h}=4LK$. Hence,
$\mathbb{E}_{\mu}[h(\widehat{T})]$ $\geqslant\mathbb{E}_{\nu}[h(T)]-L_{h}W(\mu,\nu).$
Plug into (C.11), and we get
$(L_{g}+L_{h})W(\mu,\nu)$ $\geqslant\mathbb{E}_{\nu}[(\widehat{f}(T)-f(T))\widehat{f}(T)-(\widehat{f}(T)-f(T))f(T)]$
$=\mathbb{E}_{\nu}\left[(\widehat{f}(T)-f(T))^{2}\right].$
Proof of Proposition 4.3.
@@
$=(t_{2}-t_{1})(\mathbb{E}_{{\mathcal{M}}}[Y]-\mathbb{E}_{\widehat{\mathcal{M}}}[\widehat{Y}])^{2}\leqslant(t_{2}-t_{1})W(P^{\mathcal{M}}(\bm{V}),P^{\widehat{\mathcal{M}}}(\widehat{\bm{V}}))^{2}.$
Now, suppose that $Y$ is a descendant of $T$. Let $X=\text{Pa}(T)$. Note that $f_{y}(x,t)=\mathbb{E}_{{\mathcal{M}}}[Y|X=x,T=t]$ is $L_{y}$-Lipschitz continuous with Lipschitz constant $L_{y}\leqslant(L+1)^{n_{V}}$. This is because $Y$ can be written as $Y=F_{0}(X,T,\bm{U}_{y})$ where $\bm{U}_{y}$ are latent variables independent of $T,X$ and $F_{0}$ is composition of $f_{i}$ in structure equations. The composition of Lipschitz functions is still Lipschitz and $F_{0}$ is a composition of at most $n_{V}$ $L$-Lipschitz functions. $F_{0}$ is $(L+1)^{n_{V}}$-Lipschitz and so is $f_{y}(x,t)=\mathbb{E}_{\bm{U}_{y}}[F_{0}(X,T,\bm{U}_{y})]$. Recall that $\nu,\mu$ are the observation distributions of $\mathcal{M}$ and $\widehat{\mathcal{M}}$ respectively in Proposition 4.3. Similarly, $\widehat{f}_{y}(x,t)=\mathbb{E}_{\widehat{\mathcal{M}}}[\widehat{Y}|X=x,T=t]$ is $L_{y}$-Lipschitz continuous. By Lemma C.3 and the overlap assumption, we have
$C_{W}W(\mu,\nu)$ $\geqslant\mathbb{E}_{\nu}\left[(f_{y}(X,T)-\widehat{f}_{y}(X,T))^{2}\right]$
$=\int(f_{y}(x,t)-\widehat{f}_{y}(x,t))^{2}\nu(dt|x)\nu(dx)$
$\geqslant\delta\int_{t_{1}}^{t_{2}}\int(f_{y}(x,t)-\widehat{f}_{y}(x,t))^{2}\nu(dx)P(dt).$
$\geqslant\delta\int_{t_{1}}^{t_{2}}\left(\int f_{y}(x,t)\nu(dx)-\int\widehat{f}_{y}(x,t)\nu(dx)\right)^{2}P(dt).$
Note that $\widehat{f}_{y}(x,t)$ is Lipschitz continuous, we have
$\Bigl{|}\int\widehat{f}_{y}(x,t)\nu(dx)-\int\widehat{f}_{y}(x,t)\mu(dx)\Bigl{|}\leqslant L_{y}W(\mu,\nu).$
@@
$\left(\int f_{y}(x,t)\nu(dx)-\int\widehat{f}_{y}(x,t)\nu(dx)\right)^{2}$ $\geqslant\frac{1}{2}\left(\int f_{y}(x,t)\nu(dx)-\int\widehat{f}_{y}(x,t)\mu(dx)\right)^{2}$
$\ \ \ \ \ \ \ \ \ \ \ \ -\left(\int\widehat{f}_{y}(x,t)\mu(dx)-\int\widehat{f}_{y}(x,t)\nu(dx)\right)^{2}$
$\geqslant\frac{1}{2}\left(\int f_{y}(x,t)\nu(dx)-\int\widehat{f}_{y}(x,t)\mu(dx)\right)^{2}-L_{y}^{2}W^{2}(\mu,\nu)$
$=\frac{1}{2}(\mathbb{E}_{\mathcal{M}}[Y(t)]-\mathbb{E}_{\widehat{\mathcal{M}}}[\widehat{Y}(t)])^{2}-L_{y}^{2}W^{2}(\mu,\nu).$
@@
$\int_{t_{1}}^{t_{2}}(\mathbb{E}_{\mathcal{M}}[Y(t)]-\mathbb{E}_{\widehat{\mathcal{M}}}[\widehat{Y}(t)])^{2}P(dt)$ $\leqslant\frac{2C_{W}}{\delta}W(\mu,\nu)+2L_{y}^{2}W^{2}(\mu,\nu)(t_{2}-t_{1}).$
Proof of Corollary 4.4.
In the proof of Theorem 4.2, we know with probability at least $1-O(n^{-2})$, (4.1) has feasible solutions and
$W(P_{n}^{\mathcal{M^{*}}}(\bm{V}),P^{\mathcal{M^{*}}}(\bm{V}))\leqslant O(\alpha_{n}),\quad W(P_{m_{n}}^{\theta_{n}}(\bm{V}),P^{\theta_{n}}(\bm{V}))\leqslant O(\alpha_{n}),$
where notations $\theta$ and $P^{\theta}$ are defined in the proof of Theorem 4.2 and $\theta_{n}$ is one of the minimizers of (4.1). By Lemma E.5, we know that
$W(P_{n}^{\mathcal{M^{*}}}(\bm{V}),P_{m_{n}}^{\theta_{n}}(\bm{V}))\leqslant S_{\lambda_{n}}(P_{n}^{\mathcal{M^{*}}}(\bm{V}),P_{m_{n}}^{\theta_{n}}(\bm{V}))+2(\log(m_{n}n)+1)\lambda_{n}\leqslant O(\alpha_{n}).$
@@
$W(P^{\mathcal{M^{*}}}(\bm{V}),P^{\theta_{n}}(\bm{V}))\leqslant W(P_{n}^{\mathcal{M^{*}}}(\bm{V}),P^{\mathcal{M^{*}}}(\bm{V}))+W(P^{\mathcal{M^{*}}}(\bm{V}),P^{\theta_{n}}(\bm{V}))+W(P_{m_{n}}^{\theta_{n}}(\bm{V}),P^{\theta_{n}}(\bm{V}))\leqslant O(\alpha_{n}).$
By Proposition 4.3, we get $|F_{n}-F_{*}|\leqslant O(\sqrt{\alpha_{n}})$. ∎
@@ Appendix D Experiments
The structure equations of the generative models are (3.3). We use three-layer feed-forward neural networks with width 128 for each $\widehat{f}_{i}$ and six-layer neural networks with width 128 for each $\widehat{g}_{j}$. We use the Augmented Lagrangian Multiplier (ALM) method to solve the optimization problems as in [3]. We run $600$ epochs and use a batch size of 2048 in each epoch. $m_{n}$ is set to be $m_{n}=n$. The "geomloss" package [16] is used to calculate the Sinkhorn distance. To impose Lipschitz regularization, we use the technique from [23] to do layer-wise normalization to the weight matrices with respect to infinity norm. The upper bound of the Lipschitz constant in each layer is set to be $8$. The $\tau$ in the Gumbel-softmax layer is set to be 00.
For the choice of Wasserstein ball radius $\alpha_{n}$, we use the subsampling technique from [11, Section 3.4]. We can take the criterion function in [11] to be $Q(\theta)=W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}^{\theta}}(\bm{V}))$ and its empirical estimation to be $Q_{n}(\theta)=W(P_{n}^{\mathcal{M}^{*}}(\bm{V}),P_{n}^{\mathcal{M}^{\theta}}(\bm{V}))$. In the first step, we minimize the Wasserstein distance $\widehat{\theta}^{*}_{n}=\arg\min_{\theta}W(P^{\mathcal{M}^{*}}_{n}(\bm{V}),P^{\mathcal{M}^{\theta}}_{n}(\bm{V}))$. Then, [11] propose to refine the radius $Q_{n}(\widehat{\theta}^{*}_{n})$ by taking the quantile of subsample $\{\sup_{\theta:Q_{n}(\widehat{\theta})\leqslant\beta Q_{n}(\widehat{\theta}^{*}_{n})}Q_{j,b}(\theta)\}$, where $Q_{j,b}$ denotes the criterion function estimated at $j$-th subsample of size $b$. However, it is time-consuming to solve this optimization problem many times. In practice, we set the radius $\alpha_{n}$ to be the $95\%$ quantile of $\{W(\bar{P}_{j}^{\mathcal{M}^{*}}\}(\bm{V}),P^{\mathcal{M}^{\widehat{\theta}^{*}_{n}}}_{n}(\bm{V}))$, where we use 50 subsamples $\bar{P}_{j}^{\mathcal{M}^{*}},j=1,\cdots,50$ from $P^{\mathcal{M}^{*}_{n}(\bm{V})}$ with size $b=\lfloor 15\sqrt{n}\rfloor$.